### v1

# Does basis set superposition error significantly affect post-CCSD(T) corrections?

Abstract
[…]
preprint: AIP/123-QED
I Introduction
Basis set superposition error (BSSE) is the error in the interaction energy between, for example, a dimer AB and its constituent monomers A and B when evaluated in a finite basis set. (At the complete basis set, CBS, limit, BSSE vanishes.) The classic remedy is the Boys-Bernardi counterpoise (CP) method,[1]
${\rm BSSE}$ $=$ $E[{\rm A}]+E[{\rm B}]-E[{\rm A(B)}]-E[{\rm B(A)}]$ (1)
$=$ $E[{\rm AB}]-E[{\rm A(B)}]-E{\rm[B(A)}]-(E[{\rm AB]}-E[{\rm A}]-E[{\rm B}])$
$=$ $D_{e}[{\rm raw}]-D_{e}[{\rm CP}]$
where $D_{e}$ denotes the dissociation energy, $E[{\rm A(B)}]$ denotes the energy of A in the presence of the basis functions on B, and so on.
Inclusion of BSSE in noncovalent interaction (NCI) studies is more or less standard operating procedure, especially in smaller and medium basis sets, as the CP corrections may be on the same order or magnitude as the interaction energies of interest.
Now it is indeed true that full counterpoise does not guarantee hewing closer to the CBS limit: as shown by Burns et al.[2] for orbital WFT calculations, and by Brauer et al.[3] for explicitly correlated F12 calculations, error compensation may take place between BSSE (which always overbinds) and IBSI (intrinsic basis set incompleteness, which almost invariably underbinds). Hence for small basis sets, complete neglect of BSSE may actually be beneficial, and for medium-size basis sets, ‘half-counterpoise’ (average of corrected and uncorrected interaction energies).
Figure 1: Illustration of effect of BSSE on basis set convergence for formic acid dimer. Reprinted from Figure 1 in Ref.[3], with Creative Commons license.
In computational thermochemistry, however — particularly for total atomization energies (TAE) which are the quantum chemical ‘cognates’ of heats of formation $\Delta H^{\circ}_{f}$ the IBSI overwhelms BSSE to such an extent that most researchers make no effort to apply BSSE corrections. Besides, in higher-accuracy computational thermochemistry protocols like ccCA[4, 5] by the Wilson group, HEAT by the Stanton group, Weizmann-$n$ by our own group,[6, 7, 8] and the FPD (Feller-Peterson-Dixon, see Refs.9, 10 and references therein) all entail some variant of basis set extrapolation — which, if it works properly, should eliminate BSSE altogether. (Indeed, we recently[11] exploited this fact to ‘reverse-engineer’ basis set extrapolations.)
Studies of noncovalent interactions, with rare exceptions (such as Refs.12, 13, 14) stick to the CCSD(T)[15, 16] ‘gold standard of quantum chemistry’ and ignore post-CCSD(T) corrections. However, in thermochemistry, especially for TAEs, it is well-known (see Refs.[17, 18] for early reports) that kJ/mol accuracy cannot be achieved without them. CCSD(T) in fact outperforms the more rigorous CCSDT owing to a well-established error compensation between higher-order triples, CCSDT--CCSD(T), a.k.a, $T_{3}-(T)$ (almost always antibonding), and connected quadruples (Q) (which are universally bonding).
As CCSDT(Q) and especially CCSDTQ on the one hand have very steep CPU time scalings of $O(n_{\rm occ.}^{4}N_{\rm virt.}^{5})$ and $O(n_{\rm occ.}^{4}N_{\rm virt.}^{6})$, respectively, and on the other hand, these higher-order corrections converge much more rapidly with the basis set than the overall correlation energy,[7] post-CCSD(T) corrections tend to be evaluated in very small basis sets. For instance, in W4 theory,[6] the $T_{3}-(T)$ correction is extrapolated from cc-pVDZ and cc-pVTZ basis sets, the (Q) term in a cc-pVTZ basis set, and $T_{4}-(Q)$ in just a cc-pVDZ basis set. (In W4lite theory, just CCSDT(Q)/cc-pVDZ is done for the quadruples.)
This then leads us to the main research question of the present paper: are such corrections materially affected by BSSE corrections? The issue was raised by a reviewer of Ref.[14], where we showed that post-CCSD(T) contributions for cohesive energies of water clusters approach 1 kcal/mol for isomers of (H2O)20.
II Computational Details
The CCSDT(Q),[19] CCSDT,[20] CCSD(T)Λ,[21, 22, 23, 24] and CCSD(T) calculations reported in this work were carried out using a combination of the MOLPRO 2024.1,[25] CFOUR,[26], and MRCC[27] electronic structure program systems, run on the CHEMFARM cluster of the Faculty of Chemistry at Weizmann. Owing to issues with inconsistent UHF solutions in the presence of ghost atoms, many of the small-molecule counterpoise data were generated using the MRCC interfaces of MOLPRO or CFOUR. For the noncovalent interactions, only closed-shell species are involved, and hence these calculations were carried out using standalone CFOUR (as memory permitted) or MRCC.
Reference geometries for the W4-11 thermochemical benchmark[28] (which is a subset of the larger and more recent W4-17 database[29]) were taken from the ESI of the W4-17 paper and used ‘as is’. Reference geometries for the A24 dataset were downloaded from the BEGDB database[30] of noncovalent interaction geometries: using a Python program written by one of us (ES), geometries for A24x8 were generated by compressing or stretching the intermonomer distances by the eight factors {0.9,0.95,1.0,1.05,1.1,1.25,1.50,2.0} from the familiar S66x8 database.[31]
For dissociation energies ($D_{e}$ of diatomic molecules, we applied the standard Boys-Bernardi definition. For the TAEe values (total atomization energies) of small polyatomics, we used the SSFC (site-site function counterpoise) generalization of Wells and Wilson[32], which effectively amounts to evaluating each atomic energy in the full molecular basis set.
III Results and Discussion
III.1 Total atomization energies
III.1.1 Initial check for diatomic molecules
[…]
(T) $(T)_{\Lambda}-(T)$ $T_{3}-(T)$ $T_{3}$ $(Q)$ T(Q)-(T) $(Q)_{\Lambda}-(Q)$
cc-pVDZ 0.286 0.008 0.043 0.329 0.004 0.045 0.000
cc-pVTZ 0.170 0.005 0.013 0.183 0.006 0.019 0.000
cc-pV{D,T}Za 0.008 0.008
cc-pVQZ 0.071 0.002 0.006 0.067 0.003 0.004 0.000
cc-pV5Z 0.034 0.001 0.006 0.028 0.002 0.004 0.000
haVDZ 0.214 0.008 0.026 0.238 0.002 0.026 0.000
haVTZ 0.077 0.003 0.008 0.070 0.004 0.006 0.000
haVQZ 0.037 0.001 0.010 0.028 0.001 0.009 0.000
(a) Extrapolation exponents from Table 5 of Ref.33
[…]
(T) $(T)_{\Lambda}-(T)$ $T_{3}-(T)$ $T_{3}$ $(Q)$ T(Q)-(T) $(Q)_{\Lambda}-(Q)$
cc-pVDZ 0.612 0.018 0.093 0.704 0.008 0.098 0.001
cc-pVTZ 0.353 0.024 0.039 0.389 0.015 0.053 0.003
w/o ClOO 0.353 0.020 0.036 0.388 0.015 0.050 0.002
cc-pV{D,T}Za 0.017 0.019
cc-pVQZ 0.143 0.028 0.017 0.139 0.006 0.004
haVDZ 0.443 0.052 0.052 0.487 0.005 0.006
haVTZ 0.171 0.024 0.017 0.168 0.009 0.001
(a) Extrapolation exponents from Table 5 of Ref.33 . ClOO excluded.
RMS (root mean square) BSSE corrections for a sample of 24 heavy-atom diatomics and 10 diatomic hydrides are given in Table 1.
First of all, unsurprisingly, the effect of BSSE on the difference between CCSDT(Q)Λ and CCSDT(Q) is less than 0.001 kcal/mol, and can be entirely neglected. For the difference between CCSD(T)Λ and CCSD(T) we find 0.008 kcal/mol for the cc-pVDZ and haVDZ basis sets, which tapers down to 0.001 kcal/mol for haVQZ and cc-pV5Z.
For connected quadruples (Q), the RMS BSSE is less than 0.01 kcal/mol RMS even with the cc-pVDZ basis set, and smaller still for cc-pVQZ (0.003) and haVQZ (0.001 kcal/mol). We can hence conclude that BSSE on connected quadruples is negligible even for the purposes of high-accuracy work.
The situation for higher-order connected triples, however, is somewhat different: for the cc-pVDZ basis set, we find 0.043 kcal/mol (i.e., 0.18 kJ/mol), which however drops to 0.026 kcal/mol when diffuse functions are added, and to 0.013 kcal/mol when we move things one notch up to cc-pVTZ. Both W4 and HEAT apply cc-pV{D,T}Z extrapolations to the higher-order triples: if we do so here (with extrapolation parameters taken from table V in Ref.33), the RMS BSSE is just 0.008 kcal/mol, which one may feel justified in neglecting. In fact, for the cc-pV{T,Q}Z basis sets used in W4.3 theory, the BSSE will be even more negligible.
Upon comparing RMS BSSE corrections for (T) and for all of T3 (i.e., the difference between CCSDT and CCSD), we note that for smaller basis sets like cc-pVDZ, cc-pVTZ, and haVDZ, there is more BSSE on T3 than on (T), but for larger basis sets, the roles are reversed.
In addition, if one considers the whole CCSDT(Q) - CCSD(T) difference, one finds partial mutual cancellation for BSSE for the larger basis sets, as the $T_{3}-(Q)$ and $(Q)$ corrections pull in opposite directions.
The bottom line for thermochemical applications appears to be that BSSE contributions are negligible for even high-accuracy work. Does this still bear out for polyatomics, or for noncovalent interactions?
III.1.2 Small polyatomics
While we would not be able to carry out cc-pV5Z, let alone haV5Z CCSDT(Q) calculations on polyatomics, Table 2 presents results with smaller basis sets for a subset of about three dozen triatomics from the W4-11 thermochemical benchmark.[28] While everything naturally becomes large in absolute numbers, the same basic tendencies are seen as for the diatomics: BSSE on (Q) is basically insignificant and on $(Q)_{\Lambda}-(Q)$ wholly so; BSSE on $T_{3}-(T)$ skirts the 0.1 kcal/mol edge for cc-pVDZ, but tapers down to 0.04 for cc-pVTZ, and with cc-pV{D,T}Z extrapolation, as practiced in W4 theory and HEAT, becomes negligible. We hence conclude that these thermochemical protocols require no modification to account for post-CCSD(T) BSSE unless one targets an accuracy that is likely unattainable with similar approaches.
[…]
III.2 Noncovalent interactions: the A24x8 subset
Counterpoise corrections data for the A24x8 dataset are summarized in Table 3.
[…]
(T) $T_{3}-(T)$ $(Q)$ (T) $T_{3}-(T)$ $(Q)$
cc-pVDZ(no d) 0.060 0.003 0.004 0.890 0.205 0.239
cc-pVDZ 0.067 0.002 0.003 0.759 0.162 0.208
cc-pVTZ 0.049 0.001 0.003 0.290 0.082 0.134
cc-pVQZ 0.027 0.002 0.127 0.080
haVDZ 0.049 0.002 0.297 0.140
haVTZ 0.016 0.003 0.074 0.147
haVQZ 0.007 0.029
haV5Z 0.003 0.014
Noncovalent interactions are very different in their behavior from atomization energies, in that for most noncovalent complexes, MP2 is already a decent to good starting point (except for pi-stacking and related) and CCSD-MP2 and (T) are commonly evaluated using relatively small basis sets (see, e.g., Refs.31, 34 and references therein).
Hence one might reasonably expect that post-CCSD(T) contributions will be proportionally much smaller. Admittedly, of course, the A24x8 systems are quite small, and hence post-CCSD(T) contributions might be somewhat less picayune in larger noncovalent complexes, especially at compressed geometries.
The RMS $\Delta$BSSE values are even tinier in absolute terms: 0.002 kcal/mol for $T_{3}-(T)$ and 0.003-0.004 for (Q). For smaller basis sets, these are nontrivial fractions of the actual $\Delta D_{e}$ contributions, and hence any post-CCSD(T) corrections obtained with very small basis sets, such as the unpolarized double zeta cc-pVDZ(no d), need to be regarded with caution.
For a more reasonable cc-pVTZ basis set, $\Delta$BSSE represents about 8% of the $\Delta$CP[$T_{3}-(T)$] and 13% of $\Delta$(Q).
IV Conclusions
In response to our research question, we can conclude the following:
  1. 1.
    For high-accuracy computational thermochemistry, particularly total atomization energies obtained at the W4 or HEAT levels, $T_{3}-(T)$ with the cc-pVDZ basis set carries a small but noticeable BSSE.
  2. 2.
    This is effectively removed, however, by the extrapolation of $T_{3}-(T)$ from cc-pV{D,T}Z basis sets.
  3. 3.
    BSSE on (Q) may be regarded as negligible in a thermochemistry context.
  4. 4.
    For noncovalent interactions, BSSE on both $T_{3}-(T)$ and $(Q)$ is insignificant as long as the basis set is not overly small.
Acknowledgement
This work was supported by the Israel Science Foundation (grant 1969/20) and by the Uriel Arnon Memorial Center for AI research into smart materials. JMLM thanks the Quantum Theory Project at the University of Florida and its head, Prof. John F. Stanton, for their hospitality, and Drs. James R. Thorpe, Peter R. Franke, Branko Ruscic, and Nisha Mehta as well as Prof. A. Daniel Boese (KFU Graz, Austria) for inspiring discussions. All calculations were carried out on the ChemFarm HPC cluster of the Weizmann Institute Faculty of Chemistry.
Data availability statement
Raw data can be obtained from the corresponding author upon reasonable request.